OpenStudy (anonymous):

Consider loga 18 = 2,890 and log 18 = 1,255. So, loga 10 is equal to...? a) 1,890 b) 2,302 c) 2,032 d) 1 Help pls!

OpenStudy (saifoo.khan):

"a" is in the base right?

OpenStudy (anonymous):

that's correct

OpenStudy (saifoo.khan):

Can we use calculators?

OpenStudy (anonymous):

Not sure

OpenStudy (saifoo.khan):

Why?

OpenStudy (anonymous):

loga 18 = 2,890 log 18 = 1,255 then a^2,890 = 18 and 10^1,255 = 18 so a^2,890 = 10^1,255 what do i do?

OpenStudy (anonymous):

loga 10^1,255 = 2,890

OpenStudy (anonymous):

what now?

OpenStudy (anonymous):

1,255 loga 10 = 2,890

OpenStudy (saifoo.khan):

Hey. Sorry i lost my connection.

OpenStudy (anonymous):

Np.

OpenStudy (saifoo.khan):

SOrry but i dont really get what you did. :(

OpenStudy (anonymous):

Well, loga 18 = 2,890 so a^2,890 = 18, right?

OpenStudy (anonymous):

Also, if log 18 = 1,255, then 10^1,255 = 18

OpenStudy (saifoo.khan):

Why u did a^2 ?

OpenStudy (anonymous):

That's log property. logb x = y means b^y = x.

OpenStudy (saifoo.khan):

Oho! commas were confusing me!

OpenStudy (saifoo.khan):

Well the way i was doing was, \[\frac{\log18}{\log a} = 2890\]

OpenStudy (anonymous):

Mmm I don't get that

OpenStudy (saifoo.khan):

let me thinkk.. .q

OpenStudy (saifoo.khan):

doint you think it's log(18) = 1.255 ? instead of 1,255?

OpenStudy (anonymous):

I guess that's right. We use , instead of .

OpenStudy (saifoo.khan):

oh man! now i have to do the whole thing again. lol. i thought that it's thousand..

OpenStudy (anonymous):

oh sorry xD I should've known that you guys use .

OpenStudy (saifoo.khan):

lol, Yeah. Mostly "." is used. you should mention it. :)

OpenStudy (anonymous):

Yep, I will keep that in mind.

OpenStudy (saifoo.khan):

Found it!! !!!!!!!!!!!!!!!!!!! :D It's B

OpenStudy (saifoo.khan):

It's a more mental math question! But it's really really nice one!

OpenStudy (anonymous):

please tell me how you got that!!

OpenStudy (saifoo.khan):

Can you wait for about 10-15mins?

OpenStudy (saifoo.khan):

i will brb.

OpenStudy (anonymous):

Sure.

OpenStudy (saifoo.khan):

Back! :D

OpenStudy (anonymous):

Wb :D

OpenStudy (saifoo.khan):

i am doing a mistake, hold on. i forgot the steps. lol. but the answer is b

OpenStudy (saifoo.khan):

Ok, Consider loga 18 = 2.890 and log 18 = 1.255. So, loga 10 is equal to...? a) 1.890 b) 2.302 c) 2.032 d) 1 First of all, using the log property, \[\frac{\log18}{ \log a} = 2.890\] Now they gave you the value for log 18. which is 1.255, So insert the value in the place of log 18 \[\frac{1.255}{\log a} = 2.890\]This can be written as, \[\log a = \frac{1.255}{2.890}\] \[\log a = 0.4343\] \[\large a = 10^{0.4343}\] a = 2.718

OpenStudy (saifoo.khan):

Now, \[\log_a 10\]\[\large \log_{2.718} 10\]You will get, 2.302

OpenStudy (saifoo.khan):

Phew..

OpenStudy (saifoo.khan):

i hope you get it?

OpenStudy (anonymous):

mmm whats the first property you used?

OpenStudy (anonymous):

Oh I got it

OpenStudy (saifoo.khan):

The first property is, \[\large \log_a b \to \frac{\log b}{\log a}\]

OpenStudy (anonymous):

Yep thx I got it

OpenStudy (saifoo.khan):

Welcome.

OpenStudy (anonymous):

Thank you very much for your help!! You're awesome

OpenStudy (saifoo.khan):

Welcome Welcome. I really liked the format of the question. Loved it infact.

OpenStudy (anonymous):

Lol I didn't. :D

OpenStudy (saifoo.khan):

Lol.

OpenStudy (anonymous):

You used a calc though, didnt you?

OpenStudy (saifoo.khan):

YES! How can i do without it? we have to find the values using logs!

OpenStudy (saifoo.khan):

I used it while finding the values of a.

OpenStudy (anonymous):

Tell that to the freak that invented this question. I wasn't suppose to use a calc

OpenStudy (saifoo.khan):

Haha, then how can we solve \(\Large 10^{0.4343}\) ?

OpenStudy (saifoo.khan):

We have to use a calc here. and at one more step. rest all we can do without a calc.

OpenStudy (anonymous):

Exactly, haha. Thx

OpenStudy (saifoo.khan):

Lol. welcome.

OpenStudy (anonymous):

I would never get that alone

OpenStudy (anonymous):

:P

OpenStudy (saifoo.khan):

I took this question as a challenge! :D

OpenStudy (anonymous):

Haha you're awesome indeed!!

OpenStudy (saifoo.khan):

Thankyou. ;) Btw which grade this question is about?

OpenStudy (anonymous):

No idea. Probably from 1st year of high school or something

OpenStudy (saifoo.khan):

Isn't it your question? o_O umm, no.. in 1st year of HS algebra is taught?

OpenStudy (anonymous):

Yep, I got this question from a list of exercises but it has nothing to do with homework or my school. Mmm... I don't really know, but I remember first learning about logs in 1st year of high school or 1 year earlier.

OpenStudy (saifoo.khan):

umm, okay.. nice.